# Dynamic Multimodal Prototype Learning in Vision-Language Models

### v1

@@ 1 Introduction
Figure 1: Observations of ambiguities in class names from the Oxford Flowers [33] and ImageNet [9]: (a) The class names “sword lily” and “blackberry lily” both refer to flowers from the Iridaceae family and share the word “lily”. (b) The class names “laptop” and “desktop computer” denote different types of computers.
@@
Recent advances in vision-language models (VLMs)  [1, 38, 19, 35, 51, 20] have shown remarkable success across a wide range of visual tasks. These foundation models, e.g., CLIP [38], are pre-trained on web-scale image-text pairs, achieving superior capabilities in addressing zero-shot classification [58, 41, 59, 60]. One key factor behind these successes is to find the optimal prototypes for each category. For example, CoOp [58] and CoCoOp [57] introduce learnable prompt vectors concatenated to textual prototypes. Another line of research, including TPT [41] and C-TPT [53], aims to tune class prompts during inference time for each testing sample, a strategy known as test-time adaptation.
@@
Despite the improvements brought by these methods, we empirically find that the ambiguities in class names result in textual prototypes struggling to differentiate the classes. As illustrated in Figure 1(a), the two different classes “sword lily” and “blackberry lily” exhibit high cosine similarity (0.67) in the CLIP space. This similarity arises not only from the common word “lily” but also because both flowers belong to the Iridaceae family. As a result, relying on the description-based prototypes makes it challenging to classify the images of “sword lily” and “blackberry lily” effectively (0.53 vs. 0.55 and 0.56 vs. 0.58). Furthermore, consider the class names “laptop” and “desktop computer” as depicted in Figure 1(b). Although they do not share any common same word, they still exhibit a high cosine similarity (0.69) because they both refer to different forms of computers. Consequently, test images of “laptop” and “desktop computer” have comparable similarities to both class names, making differentiation challenging (0.67 vs. 0.63 and 0.65 vs. 0.65). To this end, recent studies aim to construct high-quality prototypes in various ways, including multiple prompts generation  [28], cache-based tuning [21], and diverse data augmentation with diffusions [11]. Unfortunately, all these models focus solely on the textual domain, overlooking the information from the visual aspect [27, 2].
@@
Figure 2: Illustration of multimodal prototype learning on the ImageNet dataset [9]. (a) Process of updating multimodal prototypes. (b) Comparisons of distribution metrics.
As discussed in previous studies [2, 49, 64, 63, 37, 61], multimodal signals, such as vision and language, provide complementary knowledge that enhances the understanding of objects and class concepts. Although these class names are semantically similar, their visual characteristics provide crucial information for reducing ambiguities. In this paper, we proposed ProtoMM, a training-free multimodal prototype learning framework for adapting VLMs at test time. Moving beyond prototype learning within the single textual domain, ProtoMM aims to construct multimodal prototypes from both the textual descriptions and image stream. Specifically, for each class, we view its prototype as a discrete distribution over the corresponding label descriptions and visual particles. The visual particles collect the label-specific image features and are updated dynamically as the testing stream flows. This enables our multimodal prototype to accumulate more and more side information, improving the prediction performance. As illustrated in Figure 2(a), it is difficult to determine whether the test image belongs to the “desktop computer” or “laptop” class at time $t_{1}$. In contrast, with the help of our multimodal prototype learning, one can easily distinguish these two classes at time $t_{5}$. Furthermore, we compare the KL divergence [18] and Maximum Mean Discrepancy (MMD) [50] between our multimodal prototypes and ground truth. The results summarized in Figure 2(b) show that as time passes, the distribution discrepancy decreases significantly, e.g., KL divergence from 18.7 to 9.5 and MMD from 0.97 to 0.29.$t_{1}$ and $t_{2}$ denote the $5000$th and $10000$th sample at the test-time scenario on the ImageNet dataset [9], respectively.
@@ 2.1 Test-time Adaptation
Test-time adaptation [41, 53, 11] is an effective technique to address the distribution shift between testing and training datasets. Its core idea is to adjust the model using a streaming of unlabeled test samples. Early works [45] improve test-time adaptation performance by fine-tuning normalization layers. Motivated by the success of VLMs [38, 19], recent studies [41, 53] propose to optimize textual prompts via entropy minimization. While these methods require high computational costs due to the gradient backpropagation. To address this, another branch of works [21, 14, 56] adapting the VLMs in a non-parametric manner has emerged. Karmanov et al. [21] designs positive and negative caches to store high-confidence samples, improving the performance by calculating the similarity between the test sample and cached samples. Han et al. [14] proposes a method named DOTA [14], estimating the distribution of test samples progressively. Zhang et al. [56] extends the cache model to a dual memory network (DMN), which also works in test-time scenarios. Although these methods avoid gradient updates, they all need logit fusion for prediction, and their performance is sensitive to fusion hyperparameters.
@@ 2.2 Optimal Transport
Optimal transport (OT) [32] is initially introduced to solve the problem of how to reduce the cost when moving several items simultaneously, and then it is applied to measure the distance between distributions in the deep learning areas [23, 5, 13, 4, 65, 26], including domain adaption [7], clustering[52, 40], image classification [62, 12, 46, 25]. For example, Lazarou et al. [23] directly aligns the distribution of two sets of feature embeddings to address the transductive few-shot task. Guo et al. [13] proposes hierarchical OT to transfer the statistics from the base classes to novel classes in distribution calibration. Chen et al. [4] aims to align the features of vision and language modalities by optimizing textual prompts via OT. Recently, AWT [65], a simple and effective method adapts VLMs in zero/few-shot scenarios, which is closely related to our work. AWT formulates the image-text distance calculation as an OT problem, and solving this problem reveals the semantic correlations in the vision-language space for prediction.
@@ 3 Method
In this section, we first briefly review the preliminary knowledge of VLMs and OT [32]. Then we introduce the details of our multimodal prototype learning framework for adapting VLMs in test-time scenarios as illustrated in Figure 3.
@@ 3.1 Preliminaries
Vision-language models (VLMs). CLIP [38] as one of the well-known VLMs, consists of a vision encoder $f(\cdot)$ and a text encoder $g(\cdot)$. Given a set of $C$ class names $\{z_{c}\}_{c=1}^{C}$, CLIP first constructs the class description using a pre-defined prompt template $\Phi$, e.g., $\Phi(z_{c})$ can be chosen as “a photo of a $z_{c}$” and then obtains the prototypes by feeding the class descriptions into $g$: $\mathbf{z}_{c}=g(\Phi(z_{c}))$. For the testing image $x_{t}$ at time $t$, let $\mathbf{x}_{t}=f(x_{t})$ denotes the extracted visual feature. The zero-shot classification for $x_{t}$ is computed as:
@@
where $\mathbf{z},\mathbf{x}\in\mathbb{R}^{d}$ share the same semantic embedding space. $\mathbf{z}_{c}$ can be viewed as the prototype of $c$-th class. To improve the quality of $\mathbf{z}_{c}$, previous studies focus on test prompt tuning of $\Phi$ or generating detailed class descriptions via large language models (LLMs) [31, 36].
Optimal transport. Optimal Transport (OT) [32] is an efficient tool to measure the distance between two distributions. Specifically, consider two discrete distributions in the feature space: $P=\sum_{i=1}^{|V|}{a}_{i}\delta_{\mathbf{v}_{i}}$ and $Q=\sum_{j=1}^{|U|}{b}_{j}\delta_{\mathbf{u}_{j}}$, where $\delta$ is the Dirac delta function and $|V|$ and $|U|$ denote the number of points in $P$ and $Q$, respectively. Here, $\mathbf{a}=[a_{1},\dots,a_{|V|}]^{\top}$ and $\mathbf{b}=[b_{1},\dots,b_{|U|}]^{\top}$ denote the probability vectors that sum to 1. Given the cost matrix $\mathbf{C}\in\mathbb{R}^{{|V|}\times{|U|}}$, whose each element measures the transport cost between $\mathbf{v}_{i}$ and $\mathbf{u}_{j}$, the OT distance between distributions $P$ and $Q$ is formulated as:
@@
[…]
@@ 3.2 Distributed Feature Construction
To mimic how humans learn a new concept through different perspectives and obtain informative features, we follow strategies from previous works [65, 24, 30] and view the image and prototype as discrete distribution at the CLIP [38] space. Specifically, for each image $x_{t}$, we employ common augmentation techniques such as random crop, flip, and resize. This augments $x_{t}$ into multiple views $\{x_{t}^{n}\}_{n=1}^{N}$, including $N-1$ augmentations and the original one. Simultaneously, we query large language models [31, 36] to generate descriptive sentences for each class name. As a result, each textual prototype $z_{c}$ is enhanced with $M-1$ additional descriptions, denoted as $\{z_{c}^{m}\}_{m=1}^{M}$. Mathematically, the resulting distributions of the image and prototype are formulated as:
@@
As mentioned before, the prototype in Eq. (4) only considers the textual domain, which may not fully cover the visual details of class $z_{c}$, thereby we introduce the concept of multimodal prototype learning and aim to integrate the visual modality into $Q_{c}$. Specifically, we expand the prototype set $\{\mathbf{z}_{c}^{m}\}_{m=1}^{M}$ by appending $S$ additional visual particles:$\{\mathbf{e}_{c}^{s}\}_{s=1}^{S}$, where $\mathbf{e}_{c}\in\mathbb{R}^{d}$. Then the multimodal prototype is formulated as (We still use $Q_{c}$ for convenience):
@@
Following previous works [21, 14], we evaluate test samples sequentially in a streaming manner and hope the visual cache learns from the history, providing valuable guidance for the following samples. Before testing, the visual cache and its importance weights must be properly initialized to complete the multimodal prototype in Eq. (5). Since images are unavailable at the start of testing, it is natural to construct all visual cache of prototype $P_{c}$ by averaging corresponding textual descriptions: $\mathbf{e}_{c}^{s}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{z}_{c}^{m}.$ To determine the importance weights of each element in $P$ and $Q$, we propose a novel calculating strategy based on the input image. For the arriving image $x_{t}$, we first calculate the important weights of its $N$ visual augments according to all prototypes:
@@ 3.3 Multimodal Prototype Learning
where $\mathbf{T}_{tc}\in\mathbb{R}^{N\times(M+S)}$, whose each element denotes the transport probability from $n$-th augmentation to $m$-th prototype point in the feature space. After solving Eq. (8) using the Sinkhorn algorithm [8], we calculate the prediction probability of image $x_{t}$ as:
@@
The core idea of the visual cache is to collect high-quality samples for multimodal prototype learning. To this end, we first filter out samples whose highest prediction probability is less than a threshold: $p(y_{t}=c|\mathbf{x}_{t})\geq\tau$. For the high-quality samples, we develop a score function and only choose the augmentations with high importance. Recalling that the transport plan measures the semantic similarity between the textual prototype and visual augmentations, the score of the augmentation of $x_{t}$ thus can be calculated as $\bm{\Theta}_{t}={\bf T}_{tc}\bm{w}_{y_{t}}\in\mathbb{R}^{N}$. We then select the top-$S$ candidates $[\mathbf{x}_{t}^{(1)},\cdots,\mathbf{x}_{t}^{(S)}]^{\top}\in\mathbb{R}^{S\times d}$ according to the decreasing value sequence $[{{\theta}}_{t}^{(1)},\cdots,\theta_{t}^{(S)}]$ in $\bm{\Theta}_{t}$. The visual cache in our multimodal prototype $Q_{c}$ is updated using these selected high-quality candidates:
@@ 4.1 Setup
Datesets. Following the previous works [21, 24], we cover 11 classification datasets including diverse range categories, generic objects (ImageNet [9], Caltech [10]), scenes (SUN [48]), textures (DTD [6]), satellite images (EuroSAT [15]), actions (UCF [42]) and fine-grained categories (Pets [34], Cars [22], Flowers [33], Food [3], Aircraft [29]). Additionally, we evaluate on 4 ImageNet distribution shifted datasets [9]: ImageNetV2 [39], ImageNet-Sketch [47], ImageNet-A [17], ImageNet-R [16].
Implementation details. The experiments are built upon the pre-trained CLIP [38] model. We adapt ResNet50 (RN50) and ViT-B/16 backbones as the visual encoder, and the textual encoder is a Transformer [44]. We adhere to the same image augmentation strategies as AWT [65], including random cropping and resizing as well as flipping. For textual augmentation, we query GPT-3.5 to generate the class descriptions following the previous works [31, 65]. The total augmentation number $M$ and $N$ are both set as 50. Top-$S$ is set to 25, and threshold $\tau=0.8$. All experiments are conducted on one NVIDIA RTX 3090 GPU.
Comparison methods. We compare our method with previous state-of-the-art (SOTA) test-time adaptation methods, including TPT [41], DiffTPT [11], TDA [21], DOTA [14], DPE [55] and AWT [65]. In these comparison methods, TDA [21], AWT [65], and DOTA [14] are designed in a training-free manner. TPT [41] and DiffTPT [11] are prompt tuning methods, and DPE [55] belongs to the adapter tuning method. The optimization process of there methods are happended during the inference stage. Additionally, we compare several prompt learning methods, such as CoOp [58] and CoCoOp [57], as well as augmentation-based methods, including CUPL [36], VisDescs [30], and MetaPrompting [31].
@@
CLIP [38] 16.11 87.26 55.89 40.37 25.79 62.77 74.82 82.97 60.85 59.48 59.81 56.92
CoOp [58] 15.12 86.53 55.32 37.29 26.20 61.55 75.59 87.00 58.15 59.05 63.33 56.83
CoCoOp [57] 14.61 87.38 56.22 38.53 28.73 65.57 76.20 88.39 59.61 57.10 62.81 57.74
CuPL [36] 19.59 89.29 57.28 48.64 38.38 65.44 76.94 84.84 62.55 58.97 61.45 60.31
VisDesc [30] 16.26 88.11 54.76 41.96 37.60 65.37 76.80 82.39 59.84 58.47 59.68 58.29
TPT [41] 17.58 87.02 58.46 40.84 28.33 62.69 74.88 84.49 61.46 60.82 60.74 57.93
DiffTPT [11] 17.60 86.89 60.71 40.72 41.04 63.53 79.21 83.40 62.72 62.67 60.80 59.93
TDA [21] 17.61 89.70 57.78 43.74 42.11 68.74 77.75 86.18 62.53 64.18 61.35 61.06
DOTA [14] 18.06 88.84 58.72 45.80 47.15 68.53 78.61 87.33 63.89 65.08 61.82 62.16
DPE [55] 19.80 90.83 59.26 50.18 41.67 67.60 77.83 85.97 64.23 61.98 63.41 62.06
AWT [65] 20.31 89.57 59.22 49.46 38.80 68.21 77.18 88.22 64.91 65.02 63.01 62.17
RN50 ProtoMM 21.21 89.94 60.10 50.53 42.35 69.95 77.31 87.92 65.74 65.34 63.76 63.10
CLIP [38] 23.22 93.55 66.11 45.04 50.42 66.99 82.86 86.92 65.63 65.16 68.34 64.93
CoOp [58] 18.47 93.70 64.51 41.92 46.39 68.71 85.30 89.14 64.15 66.55 71.51 64.57
CoCoOp [57] 22.29 93.79 64.90 45.45 39.23 70.85 83.97 90.46 66.89 68.44 71.02 65.20
CuPL [36] 24.90 92.98 65.29 44.56 47.84 71.30 86.11 89.13 62.59 66.83 69.62 65.55
Visdesc [30] 24.30 94.60 64.08 44.98 54.84 70.85 85.05 88.85 67.99 67.12 68.55 66.47
TPT [41] 24.78 94.16 66.87 47.75 42.44 68.98 84.67 87.79 65.50 68.04 68.98 65.45
DiffTPT [11] 25.60 92.49 67.01 47.00 43.13 70.10 87.23 88.22 65.74 62.67 70.30 65.40
TDA [21] 23.91 94.24 67.28 47.40 58.00 71.42 86.14 88.63 67.62 70.66 69.51 67.71
DOTA [14] 25.59 94.32 69.48 47.87 57.65 74.67 87.02 91.69 69.70 72.06 70.68 69.15
DPE [55] 28.95 94.81 67.31 54.20 55.79 75.07 86.17 91.14 70.07 70.44 71.91 69.62
AWT [65] 29.22 95.40 69.80 55.56 58.40 75.07 85.54 92.23 70.00 70.70 71.26 70.28
ViT-B/16 ProtoMM 31.02 95.70 69.92 56.38 56.11 77.40 85.89 91.90 70.78 71.76 72.01 70.70
[…]

### v2

@@ 1 Introduction
Figure 1: Observations of ambiguities in class names from the Oxford Flowers [Flower] and ImageNet [ImageNet]: (a) The class names “sword lily” and “blackberry lily” both refer to flowers from the Iridaceae family and share the word “lily”. (b) The class names “laptop” and “desktop computer” denote different types of computers.
@@
Recent advances in vision-language models (VLMs)  [Flamingo, CLIP, align, LMM-R1, YangWYC023, Mimic] have shown remarkable success across a wide range of visual tasks. These foundation models, e.g., CLIP [CLIP], are pre-trained on web-scale image-text pairs, achieving superior capabilities in addressing zero-shot classification [coop, TPT, prograd, zhu2025robust]. One key factor behind these successes is to find the optimal prototypes for each category. For example, CoOp [coop] and CoCoOp [cocoop] introduce learnable prompt vectors concatenated to textual prototypes. Another line of research, including TPT [TPT] and C-TPT [C-TPT], aims to tune class prompts during inference time for each testing sample, a strategy known as test-time adaptation.
@@
Despite the improvements brought by these methods, we empirically find that the ambiguities in class names result in textual prototypes struggling to differentiate the classes. As illustrated in Figure 1(a), the two different classes “sword lily” and “blackberry lily” exhibit high cosine similarity (0.67) in the CLIP space. This similarity arises not only from the common word “lily” but also because both flowers belong to the Iridaceae family. As a result, relying on the description-based prototypes makes it challenging to classify the images of “sword lily” and “blackberry lily” effectively (0.53 vs. 0.55 and 0.56 vs. 0.58). Furthermore, consider the class names “laptop” and “desktop computer” as depicted in Figure 1(b). Although they do not share any common same word, they still exhibit a high cosine similarity (0.69) because they both refer to different forms of computers. Consequently, test images of “laptop” and “desktop computer” have comparable similarities to both class names, making differentiation challenging (0.67 vs. 0.63 and 0.65 vs. 0.65). To this end, recent studies aim to construct high-quality prototypes in various ways, including multiple prompts generation  [liu2023patch], cache-based tuning [TDA], and diverse data augmentation with diffusions [DiffTPT]. Unfortunately, all these models focus solely on the textual domain, overlooking the information from the visual aspect [LinYKPR23, multimodal_SS].
@@
Figure 2: Illustration of multimodal prototype learning on the ImageNet dataset [ImageNet]. (a) Process of updating multimodal prototypes. (b) Comparisons of distribution metrics.
[…]
@@ 2.1 Test-time Adaptation
Test-time adaptation [TPT, C-TPT, DiffTPT] is an effective technique to address the distribution shift between testing and training datasets. Its core idea is to adjust the model using a streaming of unlabeled test samples. Early works [Tent] improve test-time adaptation performance by fine-tuning normalization layers. Motivated by the success of VLMs [CLIP, align], recent studies [TPT, C-TPT] propose to optimize textual prompts via entropy minimization. While these methods require high computational costs due to the gradient backpropagation. To address this, another branch of works [TDA, dota, DMN] adapting the VLMs in a non-parametric manner has emerged. Karmanov et al. [TDA] designs positive and negative caches to store high-confidence samples, improving the performance by calculating the similarity between the test sample and cached samples. Han et al. [dota] proposes a method named DOTA [dota], estimating the distribution of test samples progressively. Zhang et al. [DMN] extends the cache model to a dual memory network (DMN), which also works in test-time scenarios. Although these methods avoid gradient updates, they all need logit fusion for prediction, and their performance is sensitive to fusion hyperparameters.
@@ 2.2 Optimal Transport
Optimal transport (OT) [monge1781memoire] is initially introduced to solve the problem of how to reduce the cost when moving several items simultaneously, and then it is applied to measure the distance between distributions in the deep learning areas [iLPC, ChenG0LC020, GuoT0ZZ22, PLOT, AWT, li25], including domain adaption [CourtyFTR17], clustering[eWWL17, ShiSY24], image classification [ease, GaoZLG23, wang2023tuning, li2023patchct]. For example, Lazarou et al. [iLPC] directly aligns the distribution of two sets of feature embeddings to address the transductive few-shot task. Guo et al. [GuoT0ZZ22] proposes hierarchical OT to transfer the statistics from the base classes to novel classes in distribution calibration. Chen et al. [PLOT] aims to align the features of vision and language modalities by optimizing textual prompts via OT. Recently, AWT [AWT], a simple and effective method adapts VLMs in zero/few-shot scenarios, which is closely related to our work. AWT formulates the image-text distance calculation as an OT problem, and solving this problem reveals the semantic correlations in the vision-language space for prediction.
@@ 3 Method
In this section, we first briefly review the preliminary knowledge of VLMs and OT [monge1781memoire]. Then we introduce the details of our multimodal prototype learning framework for adapting VLMs in test-time scenarios as illustrated in Figure 3.
@@ 3.1 Preliminaries
Vision-language models (VLMs). CLIP [CLIP] as one of the well-known VLMs, consists of a vision encoder $f(\cdot)$ and a text encoder $g(\cdot)$. Given a set of $C$ class names $\{z_{c}\}_{c=1}^{C}$, CLIP first constructs the class description using a pre-defined prompt template $\Phi$, e.g., $\Phi(z_{c})$ can be chosen as “a photo of a $z_{c}$” and then obtains the prototypes by feeding the class descriptions into $g$: $\mathbf{z}_{c}=g(\Phi(z_{c}))$. For the testing image $x_{t}$ at time $t$, let $\mathbf{x}_{t}=f(x_{t})$ denotes the extracted visual feature. The zero-shot classification for $x_{t}$ is computed as:
@@
where $\mathbf{z},\mathbf{x}\in\mathbb{R}^{d}$ share the same semantic embedding space. $\mathbf{z}_{c}$ can be viewed as the prototype of $c$-th class. To improve the quality of $\mathbf{z}_{c}$, previous studies focus on test prompt tuning of $\Phi$ or generating detailed class descriptions via large language models (LLMs) [metaprompt, CuPL].
Optimal transport. Optimal Transport (OT) [monge1781memoire] is an efficient tool to measure the distance between two distributions. Specifically, consider two discrete distributions in the feature space: $P=\sum_{i=1}^{|V|}{a}_{i}\delta_{\mathbf{v}_{i}}$ and $Q=\sum_{j=1}^{|U|}{b}_{j}\delta_{\mathbf{u}_{j}}$, where $\delta$ is the Dirac delta function and $|V|$ and $|U|$ denote the number of points in $P$ and $Q$, respectively. Here, $\mathbf{a}=[a_{1},\dots,a_{|V|}]^{\top}$ and $\mathbf{b}=[b_{1},\dots,b_{|U|}]^{\top}$ denote the probability vectors that sum to 1. Given the cost matrix $\mathbf{C}\in\mathbb{R}^{{|V|}\times{|U|}}$, whose each element measures the transport cost between $\mathbf{v}_{i}$ and $\mathbf{u}_{j}$, the OT distance between distributions $P$ and $Q$ is formulated as:
@@
where $\mathbf{T}\in\mathbb{R}^{|V|\times|U|}$ is the transport plan to be learned, which indicates the transport probability from source point in $P$ to target point in $Q$. $\langle\cdot,\cdot\rangle$ denotes the Frobenius dot-product and $\mathbbm{1}_{|V|}$ refers to $|V|$-dimensional all-one vector. However, directly computing Eq. (2) is often computationally intensive. Following previous work [iLPC, PLOT], we employ the Sinkhorn algorithm [Cuturi13], which introduces an entropic regularization term to facilitate efficient optimization. The resulting objective is then given by:
@@ 3.2 Distributed Feature Construction
To mimic how humans learn a new concept through different perspectives and obtain informative features, we follow strategies from previous works [AWT, WAC, VisDesc] and view the image and prototype as discrete distribution at the CLIP [CLIP] space. Specifically, for each image $x_{t}$, we employ common augmentation techniques such as random crop, flip, and resize. This augments $x_{t}$ into multiple views $\{x_{t}^{n}\}_{n=1}^{N}$, including N1N-1 augmentations and the original one. Simultaneously, we query large language models [metaprompt, CuPL] to generate descriptive sentences for each class name. As a result, each textual prototype $z_{c}$ is enhanced with M1M-1 additional descriptions, denoted as $\{z_{c}^{m}\}_{m=1}^{M}$. Mathematically, the resulting distributions of the image and prototype are formulated as:
@@
[…]
@@
Following previous works [TDA, dota], we evaluate test samples sequentially in a streaming manner and hope the visual cache learns from the history, providing valuable guidance for the following samples. Before testing, the visual cache and its importance weights must be properly initialized to complete the multimodal prototype in Eq. (5). Since images are unavailable at the start of testing, it is natural to construct all visual cache of prototype $P_{c}$ by averaging corresponding textual descriptions: $\mathbf{e}_{c}^{s}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{z}_{c}^{m}.$ To determine the importance weights of each element in $P$ and $Q$, we propose a novel calculating strategy based on the input image. For the arriving image $x_{t}$, we first calculate the important weights of its $N$ visual augments according to all prototypes:
@@ 3.3 Multimodal Prototype Learning
where $\mathbf{T}_{tc}\in\mathbb{R}^{N\times(M+S)}$, whose each element denotes the transport probability from $n$-th augmentation to $m$-th prototype point in the feature space. After solving Eq. (8) using the Sinkhorn algorithm [Cuturi13], we calculate the prediction probability of image $x_{t}$ as:
@@
The core idea of the visual cache is to collect high-quality samples for multimodal prototype learning. To this end, we first filter out samples whose highest prediction probability is less than a threshold: $p(y_{t}=c|\mathbf{x}_{t})\geq\tau$. For the high-quality samples, we develop a score function and only choose the augmentations with high importance. Recalling that the transport plan measures the semantic similarity between the textual prototype and visual augmentations, the score of the augmentation of $x_{t}$ thus can be calculated as $\bm{\Theta}_{t}={\bf T}_{tc}\bm{w}_{y_{t}}\in\mathbb{R}^{N}$. We then select the top-$S$ candidates $\mathbf{x}_{t}^{(1)},\cdots,\mathbf{x}_{t}^{(S)}]^{\top}\in\mathbb{R}^{S\times d}$ according to the decreasing value sequence ${{\theta}}_{t}^{(1)},\cdots,\theta_{t}^{(S)}]$ in $\bm{\Theta}_{t}$. The visual cache in our multimodal prototype $Q_{c}$ is updated using these selected high-quality candidates:
@@ 4.1 Setup
Datesets. Following the previous works [TDA, WAC], we cover 11 classification datasets including diverse range categories, generic objects (ImageNet [ImageNet], Caltech [Caltech]), scenes (SUN [SUN]), textures (DTD [DTD]), satellite images (EuroSAT [Eurosat]), actions (UCF [UCF101]) and fine-grained categories (Pets [OxfordPet], Cars [Cars], Flowers [Flower], Food [Food-101], Aircraft [FGVC]). Additionally, we evaluate on 4 ImageNet distribution shifted datasets [ImageNet]: ImageNetV2 [ImageNetV2], ImageNet-Sketch [ImageNetSketch], ImageNet-A [ImageNetA], ImageNet-R [ImageNetR].
Implementation details. The experiments are built upon the pre-trained CLIP [CLIP] model. We adapt ResNet50 (RN50) and ViT-B/16 backbones as the visual encoder, and the textual encoder is a Transformer [transformer]. We adhere to the same image augmentation strategies as AWT [AWT], including random cropping and resizing as well as flipping. For textual augmentation, we query GPT-3.5 to generate the class descriptions following the previous works [metaprompt, AWT]. The total augmentation number $M$ and $N$ are both set as 50. Top-$S$ is set to 25, and threshold $\tau=0.8$. All experiments are conducted on one NVIDIA RTX 3090 GPU.
Comparison methods. We compare our method with previous state-of-the-art (SOTA) test-time adaptation methods, including TPT [TPT], DiffTPT [DiffTPT], TDA [TDA], DOTA [dota], DPE [DPE] and AWT [AWT]. In these comparison methods, TDA [TDA], AWT [AWT], and DOTA [dota] are designed in a training-free manner. TPT [TPT] and DiffTPT [DiffTPT] are prompt tuning methods, and DPE [DPE] belongs to the adapter tuning method. The optimization process of there methods are happended during the inference stage. Additionally, we compare several prompt learning methods, such as CoOp [coop] and CoCoOp [cocoop], as well as augmentation-based methods, including CUPL [CuPL], VisDescs [VisDesc], and MetaPrompting [metaprompt].
@@
CLIP [CLIP] 16.11 87.26 55.89 40.37 25.79 62.77 74.82 82.97 60.85 59.48 59.81 56.92
CoOp [coop] 15.12 86.53 55.32 37.29 26.20 61.55 75.59 87.00 58.15 59.05 63.33 56.83
CoCoOp [cocoop] 14.61 87.38 56.22 38.53 28.73 65.57 76.20 88.39 59.61 57.10 62.81 57.74
CuPL [CuPL] 19.59 89.29 57.28 48.64 38.38 65.44 76.94 84.84 62.55 58.97 61.45 60.31
VisDesc [VisDesc] 16.26 88.11 54.76 41.96 37.60 65.37 76.80 82.39 59.84 58.47 59.68 58.29
TPT [TPT] 17.58 87.02 58.46 40.84 28.33 62.69 74.88 84.49 61.46 60.82 60.74 57.93
DiffTPT [DiffTPT] 17.60 86.89 60.71 40.72 41.04 63.53 79.21 83.40 62.72 62.67 60.80 59.93
TDA [TDA] 17.61 89.70 57.78 43.74 42.11 68.74 77.75 86.18 62.53 64.18 61.35 61.06
DOTA [dota] 18.06 88.84 58.72 45.80 47.15 68.53 78.61 87.33 63.89 65.08 61.82 62.16
DPE [DPE] 19.80 90.83 59.26 50.18 41.67 67.60 77.83 85.97 64.23 61.98 63.41 62.06
AWT [AWT] 20.31 89.57 59.22 49.46 38.80 68.21 77.18 88.22 64.91 65.02 63.01 62.17
RN50
ProtoMM 21.21 89.94 60.10 50.53 42.35 69.95 77.31 87.92 65.74 65.34 63.76 63.10
CLIP [CLIP] 23.22 93.55 66.11 45.04 50.42 66.99 82.86 86.92 65.63 65.16 68.34 64.93
CoOp [coop] 18.47 93.70 64.51 41.92 46.39 68.71 85.30 89.14 64.15 66.55 71.51 64.57
CoCoOp [cocoop] 22.29 93.79 64.90 45.45 39.23 70.85 83.97 90.46 66.89 68.44 71.02 65.20
CuPL [CuPL] 24.90 92.98 65.29 44.56 47.84 71.30 86.11 89.13 62.59 66.83 69.62 65.55
Visdesc [VisDesc] 24.30 94.60 64.08 44.98 54.84 70.85 85.05 88.85 67.99 67.12 68.55 66.47
TPT [TPT] 24.78 94.16 66.87 47.75 42.44 68.98 84.67 87.79 65.50 68.04 68.98 65.45
DiffTPT [DiffTPT] 25.60 92.49 67.01 47.00 43.13 70.10 87.23 88.22 65.74 62.67 70.30 65.40
TDA [TDA] 23.91 94.24 67.28 47.40 58.00 71.42 86.14 88.63 67.62 70.66 69.51 67.71
DOTA [dota] 25.59 94.32 69.48 47.87 57.65 74.67 87.02 91.69 69.70 72.06 70.68 69.15
DPE [DPE] 28.95 94.81 67.31 54.20 55.79 75.07 86.17 91.14 70.07 70.44 71.91 69.62
AWT [AWT] 29.22 95.40 69.80 55.56 58.40 75.07 85.54 92.23 70.00 70.70 71.26 70.28
ViT-B/16
ProtoMM 31.02 95.70 69.92 56.38 56.11 77.40 85.89 91.90 70.78 71.76 72.01 70.70
Table 1: Performance comparisons of accuracy ($\%$) on the cross-domain benchmark. All the compared methods are built upon CLIP-RN50 or CLIP-ViT-B/16 backbone. The evaluation metric Average is calculated by taking the mean accuracy across all 11 datasets. The best and second best results are denoted in bold and underline, respectively. refers to our implementation using their publicly released code.